OpenStudy (anonymous):

Transform the following polar equation into an equation in rectangular coordinates:? r=16 Choices: A. (x-4)^2+y^2=4 B.x=4 C. xy=16 D. x^2+y^2=4 I could really use help! thanks :)

OpenStudy (anonymous):

Sounds like a circle

OpenStudy (amistre64):

r = 16 is a polar equation where the radius is always a distance of 16 from the center of a circular type thingamabob

OpenStudy (amistre64):

b and c are out; and there is no indication that we moved from origin so my guess is d

OpenStudy (anonymous):

thank you!

OpenStudy (amistre64):



OpenStudy (amistre64):

also; r^2 = x^2 + y^2 which really almost looks like d

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=polar+r%3D16 which should amount to: x^2 + y^2 = 256